If f(x) is an odd periodic function with period 2 then f(4) equals
  1. -4
  2. 4
  3. 2
  4. 0

Grade:Select Grade

4 Answers

anuj
8 years ago
zero
Rakhi Agrawal
8 years ago
please explain
Shubham Singh
8 years ago
f(x) is odd means f(-x) = -f(x)
so, f(0) = -f(0)
or, f(0) = 0
its period is 2 so f(0) = f(2) = f(4) = 0
ankit singh
3 years ago
function is odd means if be put x instead of -x the funcf(-x)=-f(x) so f(0)=-f(0) or f(0)=0,its priod is 2 so f(2)=f(4)=0tion become neagative from  the function
this the answer by forum expert faculty
ankit singh
nit bhopal